HUBBLE SPACE TELESCOPE DAILY REPORT #2461 PERIOD COVERED: 0000Z (UTC) 09/20/99 - 0000Z (UTC) 09/21/99 Daily Status Report as of 264/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8172 (STIS Deep DLA Imaging Survey II) The Space Telescope Imaging Spectrograph (CCD) was used to make deep observations of quasar/alpha galaxy PKS0201+11. The HST imaging, together with ground-based spectroscopy, will unify absorption-line studies and faint galaxy imaging studies, providing a comprehensive picture of the history of assembly, star formation, gas consumption, and chemical enrichment in the Damped Lyman-Alpha galaxies. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 8436 (POMS Test Proposal: WFII Backup Parallel Archive Proposal II) The WF/PC-2 was used to make archive observations parallel to the prime science observations as part of a POMS test proposal. This is a POMS test proposal designed to simulate scientific plans. The observations were completed as planned, and no anomalies were reported. 1.3 Completed Two Sets of WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was used to make a ultraviolet Earth flat calibration image. This is to improve the quality of pipeline the ultraviolet flatfields. As described in HSTAR 7261 and yesterday’s report, the acquisition/re-acquisitions for the second iteration of this proposal failed and the take data flag remained down. Otherwise, the observations were completed as scheduled, and no problems were reported. 1.4 Completed Two Sets of WF/PC-2 8345 (Multipolar Bubbles and Jets in Very Low Excitation Planetary Nebulae - A WF/PC-2 H-Alpha Imaging Survey) The WF/PC-2 was used to make H-Alpha observations of the very low excitation planetary nebula PK275-03D1 in order to study its multipolar bubbles and jets. Comparison of the morphology of this young PNe sample with more evolved PNe populations will set new constraints on models for the production of jet-like outflows and nebular evolution. The observations were completed as planned, and no problems were reported. 1.5 Completed Three Sets of STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.6 Completed Five Sets of WF/PC-2 8442 (WF/PC-2 Cycle 8 Standard Darks) The WF/PC-2 was used to make a series of Cycle 8 standard dark calibration observations. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. As described in HSTAR 7261 and yesterday’s report, the acquisition/re-acquisitions for the third iteration of this proposal failed and the take data flag remained down. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.7 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.8 Completed Two Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.9 Completed FGS 8309 (A Link Between Massive Binary Stars and Non-thermal Radio Emission) The Fine Guidance Sensors were used to make astrometric observations of the massive Wolf-Rayet binary star WR105 and several nearby reference stars. Non-thermal radio emission in Wolf-Rayet stars is explained in terms of synchrotron emission from shocks in the stellar wind. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Two Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Two Sets of WF/PC-2 6603 (Probing Proto-Planetary Disks in the Orion Nebula) The WF/PC-2 was used to make images of the proto-planetary disk NGC-1976 which is located in the Orion Nebula. This is to study the fundamental physical properties and structures of recently discovered, and uniquely visible, proto-planetary disks around young stars. As described in HSTAR 7261 and yesterday’s report, the acquisition/re-acquisitions for the first iteration of this proposal failed and the take data flag remained down. Then, during the second iteration, the high voltage was unexpectedly turned off as described in 2.3 and HSTAR 7264. This occurred about 45 minutes before this eight-and-one-half-hour proposal had completed. No further problems were reported. 1.12 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was to be used to make a CCD bias calibration observation. As described in HSTAR 7261 and yesterday’s report, the acquisition/re-acquisitions for this proposal failed and the take data flag remained down. 1.13 Completed STIS/MA1 8431 (Doppler Test of Far UV MAMA with Jenkins Slit) The Space Telescope Imaging Spectrograph (MA1) was to be used to execute a Doppler test to measure the limiting resolving power of the far-ultraviolet MAMA using the Jenkins 0.100x0.030 slit. Observations were to be made of standard star HD141556. About one hour into this approximate five-hour proposal, the event described in 2.3 and HSTAR 7267 occurred, causing the remaining observations to be lost. Otherwise, the observations were completed as planned, and no further anomalies were noted. 1.14 Completed STIS/CCD/MA1 8188 (Snapshot Spectroscopic Diagnostic Survey of the Symbiotic Stars) The Space Telescope imaging Spectrograph (CCD and MA1) was used to make ultraviolet observations of star V835-Cen as part of a snapshot spectroscopic diagnostic survey of symbiotic stars. Using these data we will be able to establish meaningful statistical limits on the nature of hot component sources of ionization and fundamental wind parameters, which are vital steps toward understanding the accretion hydrodynamics and the evolutionary status of symbiotic systems. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 6 Successful: 6 Scheduled Re-acquisitions: 10 Successful: 6 Yesterday’s report presented HSTAR 7261, describing acquisition/re-acquisition failures due to the search radius limit being exceeded. Since that report, additional re-acquisition failures at 263/1043Z, 263/1129Z, 263/1221Z, and 263/1358Z occurred from the same event. 2.2 FHST Updates: Scheduled: 11 Successful: 11 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was reset two times. At 263/154441Z, several STIS MAMA-2 status buffer messages were received indicating that the high voltage had been turned off. HSTAR 7264 was written. The proposal detailed in 1.11 was affected. Per HSTAR 7267, at 263/194915Z, an NSSC-1 status buffer message was generated for a CU/SDF error, along with a host of STIS error messages. Analysis by SEs determined that this was caused by a command overlap in the Command Timeline. The proposal outlined in 1.13 was affected. Per ROP NS-3, the NSSC-1 status buffer was dumped and reset at 263/2229Z. Then, using ROP NS-12, the STIS flight software error counter was reset at 263/2232Z. Finally, as directed by ROP NS-5, two SI C&DH errors were reset at 263/2246Z. A TTR was written for a period of no GCMR capability, beginning at 264/0011Z, due to an NCC software problem. Another TTR was written (for the same problem and reason) for a period beginning at 264/0247Z. Finally, a TTR was written for another period of no GCMR capability that resulted in the non-recoverable loss of 4k engineering data for the period 264/043100Z until 264/043400Z. At 264/031250Z, HSTAR 7268 tells about a STIS status buffer message was received indicating that the MAMA-2 low voltage was not allowed to turn-on due to an event flag that had been set by the flight software. This was determined to be related to the problems related to HSTAR 7264 (described above). STIS/CCD proposal 8238, occurring during the period of the next report, was affected. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH